OpenStudy (bekkah323):

Multiply and express your results using scientific notation? (3.1*10^5)(4.5*10^-3)

OpenStudy (anonymous):

Are you allowed to use a calculator? My calculator got 1,395 and in Scientific notation it would be 1.4x10^3.

OpenStudy (bekkah323):

Yes, and how did you get that

OpenStudy (anonymous):

multiply \(3.1\times 4.5\) then also multiply \(10^5\times 10^{-3}\) the first is just a multiplication, the second you do by \(10^5\times 10^{-3}=10^{5-3}=10^2\)

OpenStudy (anonymous):

since \(3.1\times 4.5=13.95\) you get \[13.95\times 10^2\] to put in scientific notation, move the decimal place left one place, and add one to the exponent you get \(1.395\times 10^3\)

OpenStudy (bekkah323):

thank you